OpenStudy (anonymous):

find the second derivative for 4x^3/2 its quite simple but i keep having problems with fractions

OpenStudy (mathstudent55):

\(\large y = 4x^{\frac{3}{2}} \) \(\large y' = \dfrac{3}{2}(4)x^{\frac{1}{2}} = 6x^{\frac{1}{2}} \) \(\large y'' = (6x^{\frac{1}{2}})' = \dfrac{1}{2}(6) {x^{-\frac{1}{2}}} = 3{x^{-\frac{1}{2}}} = \dfrac{3}{\sqrt{x}} \)